A locomotive with a mass of 3E5 is accelerating at 0.4m/s2 up an incline of 1 in 20. The rolling resistance is 20 kN of the locomotive. At the instant when the speed of the locomotive is 63 km/h, calculate:

6.1. The power required. (10)

6.2. The kinetic energy of the locomotive. (8)


1
Expert's answer
2021-08-04T04:40:26-0400

"m = 3*10^5kg"

"a = 0.4\frac{m}{s^2}"

"F_{rol} = 20000N"

"v=63\frac{km}{h }=17.5\frac{m}{s}"

"P = F_{loc}v"

"F = ma"

"F = F_{loc}-F_{rol}-mg\sin\alpha"

"\sin \alpha= \frac{1}{20}=0.05"

"F_{loc}=ma+F_{rol}+mg\sin\alpha=287000N"

6.1

"P = F_{loc}v= 5022500W"

6.2

"E_k = \frac{mv^2}{2}=45937500J"


"\text {Answer}:"

6.1 P = 5022500W

6.2 Ek = 45937500J